Four years from now, Malia will be 3 times as old as Allison is then. The sum of their ages is 2 more than 5 times Allison's age now.
What are the starting equations?

Answer by ikleyn(52776)

.
Four years from now, Malia will be 3 times as old as Allison is then. The sum of their ages is 2 more than 5 times Allison's age now.
What are the starting equations?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

The starting equations are

M + 4 = 3*(A+4),    (1)   ("Four years from now, Malia will be 3 times as old as Allison is then.")

M + A = 2 + 5A.     (2)   ("The sum of their ages is 2 more than 5 times Allison's age now.")

I do not solve them, because you don't ask about it.

Good luck !

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.